OpenStudy (anonymous):

What is the value of the expression? (–4)4 have to work it out and not sure how to do it i use to know how but i forgot lol

OpenStudy (anonymous):

@Ajax

OpenStudy (sidsiddhartha):

\[a*b=ab\\-a*b=-ab\\a*-b=-ab\\-a*-b=ab\]

OpenStudy (sidsiddhartha):

can u do it now?

OpenStudy (anonymous):

i stilll don't understand it.

OpenStudy (sidsiddhartha):

\[-4*4=-(4*4)\] now (4*4)=?

OpenStudy (anonymous):

u mean 4x4?

OpenStudy (sidsiddhartha):

yeah

OpenStudy (anonymous):

ok thanks

OpenStudy (sidsiddhartha):

okey so,whats the answer u got?